OpenStudy (anonymous):

Check my work? 7x - 3y = 20 5x + 3y = 16 I added them together 12x + 0 = 36 12x = 36 Divided by 12 x = 3 Substitute 3 into either equation. 7(3) - 3y = 20 21 - 3y - 21 = 20 - 21 -3y = -1 Divide each side by -3 y = -1/3 (3, -1/3)?

OpenStudy (anonymous):

Looks good except the y value is not negative. It should look something like (3, 1/3)

OpenStudy (anonymous):

Okay, thank you!

OpenStudy (anonymous):

Yep

OpenStudy (eyust707):

7(3) - 3y = 20 21 - 3y - 21 = 20 - 21 -3y = -1 all good.. then y = -1/-3 y= 1/3

OpenStudy (anonymous):

Wait do either of you know what method this is called?